OpenStudy (anonymous):

One canned juice drink is 25% orange juice, another is 5% orange juice. How many liters of each should be mixed together to get a 20L that is 16% orange juice? How many liters of 25% orange juice should be in the mixture? How many liters of 5% orange juice should be in the mixture?

OpenStudy (ganpat):

let x be juice containing 25% orange and y be juice containing 5 % orange.. so you have two equations now.. x + y = 20 and, 25x + 5y = 16 (20) can u solve further ??

OpenStudy (ganpat):

multiply eqnt 1 by 5.. 5x + 5y = 100 subtract equatn 2.. 20x = 120 x= 6 liters... and y = 14 liters...